### v1

# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

Abstract
Existing point cloud representation learning tend to learning the geometric distribution of objects through data-driven approaches, emphasizing structural features while overlooking the relationship between the local information and the whole structure. Local features reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo elastic deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric properties. Inspired by this, we propose a physics-driven self-supervised learning method for point cloud representation, which captures the relationship between parts and the whole by constructing a local-whole force propagation mechanism. Specifically, we employ a dual-task encoder-decoder framework, integrating the geometric modeling capability of implicit fields with physics-driven elastic deformation. The encoder extracts features from the point cloud and its tetrahedral mesh representation, capturing both geometric and physical properties. These features are then fed into two decoders: one learns the whole geometric shape of the point cloud through an implicit field, while the other predicts local deformations using two specifically designed physics information loss functions, modeling the deformation relationship between local and whole shapes. Experimental results show that our method outperforms existing approaches in object classification, few-shot learning, and segmentation, demonstrating its effectiveness.
1 Introduction
Point clouds, as a common data structure, provide a compact representation of 3D objects, capable of precisely describing the geometric shape and spatial distribution of objects. Due to their rich geometric information and broad range of applications, point clouds play a crucial role in 3D vision tasks, such as 3D shape analysis [52], object detection [38], and segmentation [15, 8]. The success of these tasks depends on the accurate modeling of point cloud features, while the irregularity and sparsity of point clouds make learning their effective representations a challenging but essential task.
Figure 1: Deformation visualization. When different magnitudes of forces $({f_{1}}<{f_{2}})$ are applied from top to bottom on the hat, it undergoes varying degrees of deformation.
Existing methods primarily rely on data-driven approaches to capture geometric distributions and structural features of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [30], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. PointNet++ [31] builds upon this by introducing a hierarchical structure to capture local features, allowing for better capture of structural information in local regions. DGCNN [45] utilizes dynamic graph convolutional networks to capture local geometric structural features. In unsupervised point cloud representation learning, generative methods [56, 55, 23, 13] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [24], generative adversarial networks (GAN) [7], or autoregressive (AR) [39, 23]. Discriminative methods [49, 16, 47, 22] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [14, 34] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
Specifically, existing methods often expand local information to the whole level through multi-scale operations (e.g., the hierarchical structure of PointNet++) or max-pooling operations (e.g., global feature aggregation in PointNet). This process of aggregating information solely in the feature space does not explicitly model and capture the interrelationship between local and whole structures. For example, DGCNN captures local geometric structures using dynamic graph convolutional networks, but its layer-by-layer expansion of local information remains a passive, lacking explicit modeling of the relationship between part and whole structural features.
In the real world, objects undergo elastic deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric properties. As shown in Figure 1, objects change to different degrees when forces of different magnitudes are applied to them. Based on this, we propose a physics-driven self-supervised learning method for point cloud representation, which captures the deformation effects between local and whole shapes by introducing the force propagation mechanism from elasticity theory. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of the point cloud from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, simulating the force propagation relationship between local and global structures. The main contributions of our method are as follows:
$\bullet$ We propose a novel self-supervised learning method that explicitly captures the relationship between fine-grained features and whole structure in point clouds by combining physical laws with geometric modeling. This is also the first method to apply elastic mechanics to self-supervised learning.
$\bullet$ We designed two key loss functions to learn the physical characteristics of objects: a data-fidelity loss and a physics-informed loss. These two loss functions work synergistically during the optimization process to enforce physical consistency in the point cloud.
$\bullet$ We demonstrate the effectiveness of our method through extensive evaluations on multiple downstream tasks, including object classification, few-shot learning, and segmentation. Additionally, the incorporation of physical information enhances the understanding of 3D point clouds.
2 Related Works
Point Cloud Representation Learning: In recent years, the advancement of deep learning techniques has enabled direct utilization of point cloud data for learning intrinsic feature representations of shapes. Early methods, such as PointNet [30], used shared MLPs and max pooling operations to learn global features of point clouds. Subsequently, PointNet++ [31] introduced a hierarchical structure and local feature learning mechanisms to address the issue of PointNet’s insensitivity to local features in point clouds. DGCNN [45] further improves local feature extraction by dynamically constructing graphs based on point neighborhoods. Similarly, recent studies [32, 33, 19] have focused on learning local feature information from point clouds, aiming to extract as many semantic features as possible. However, the methods often treat point clouds as static geometric distributions, overlooking the explicit relationship between local and whole structures. This limitation, especially in the case of object deformation, makes it difficult to capture the impact of local changes on the whole structure.
Figure 2: The overall architecture of the proposed method is illustrated. It consists of two modules: the implicit feature learning module, which represents the point cloud as a continuous surface, helping the network better learn the whole shape features of the object; and the physics information awareness module, which learns the relationship between local features and whole geometry through the elastic deformation of the object, enhancing the model’s ability to learn local details.
Unsupervised Point Cloud Representation Learning: Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [44, 56, 20, 50, 13, 41]. For example, HC-VAE [20] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [50] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [55, 23, 56, 53] has also become a prominent research focus. MaskSurf [55] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [56] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process. However, generative methods tend to learn the latent representations of point clouds through whole reconstruction tasks, overlooking local features and their impact on the whole structure.
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [36, 37, 49, 46, 26, 4]. Jigsaw3D [36] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [29] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [16, 21, 49] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [49] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [16] is an extension of the method [11] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [22] guides network learning by maximizing the mutual information between two point clouds. However, discriminative methods typically rely on a large number of sample pairs and carefullydesigned learning strategies, making it challenging to effectively learn the interactions between local deformations and global shapes.
In addition, some researchers have focused on incorporating cross-model [1, 47, 54, 18, 28] methods into discriminative methods. CrossPoint [1] integrates both 2D and 3D modal information, generating more robust point cloud features. Building upon CrossPoint, CrossNet [47] further divides the images into RGB and grayscale for comparative analysis. TCMSS [12] parameterizes point cloud and image data as Dirichlet distributions and then aligns the data from different modalities.
3 Methods
3.1 Overview
The framework of our proposed method is illustrated in Figure 2. It is primarily composed of two modules: The first module is the implicit feature learning module, where the point cloud is input into an autoencoder to learn its implicit features. The implicit field represents the point cloud as a continuous surface, and by leveraging the distance function from a query point to the nearest surface, it can uniformly handle objects of arbitrary shapes, helping the network better learn the whole shape features of the object. The second module is the physical information awareness module, which captures the mechanical characteristics of local deformations in the point cloud to model the relationship between local features and the whole geometric shape. Our autoencoder-based dual-task learning framework not only provides a more comprehensive representation of geometric structures but also incorporates physical constraints, enabling more efficient learning of point cloud representations with real-world significance.
3.2 Implicit Feature Learning
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${{\Phi}_{\theta}}$ and decoder ${{\Psi}_{\eta}}$ to learn the implicit features of the point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$. The encoder ${{\Phi}_{\theta}}$ takes the ${P_{i}}$ as input and maps it to an $m$-dimensional latent vecoter ${z_{\eta}}$, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$. Where we use the network backbones of PointNet [30] and DGCNN [45] as ${{\Phi}_{\theta}}$ to learn the point cloud features, respectively. Subsequently, we concatenate latent vector ${z_{\eta}}$ with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Then, the decoder ${{\Psi}_{\eta}}$ takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud ${P_{i}}$. The decoder ${{\Psi}_{\eta}}$ consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
Our training objective is to minimize the error between the distance from the query point $q$ to the implicit surface of the point cloud ${P}_{i}$, as predicted by the implicit function $\delta_{i}({z_{\eta}^{q}})$, and the ground truth distance ${\delta}_{gt}(q)$:
${\theta^{*}},{\eta^{*}}=\mathop{\arg\min}\limits_{\theta,\eta}\left({\delta_{i}(z_{\eta}^{q})),{\delta_{gt}}\left(q\right)}\right)$ (1)
where ${\delta}_{gt}(q)$ represents the nearest neighbor distance from the query point $q$ to the points on the point cloud ${{P}_{i}}$, expressed using the unsigned distance function value [6]. Therefore, the implicit loss ${\cal L}_{imp}$ is defined as:
${{\cal L}_{imp}}=\frac{1}{{|\mathcal{D}|K}}\sum\limits_{i=1}^{|\mathcal{D}|}{\sum\limits_{q\in Q}^{K}{\left|{\left|{{\delta_{i}}(z_{\eta}^{q})}\right|-{\delta_{gt}(q)}}\right|}}$ (2)
where $K$ is the number of the query point $q$ randomly sampled within the unit space, and $Q$ is the set of query points.
3.3 Physical Information Awareness
For a point cloud ${{P}_{i}}$, we first use 3D Delaunay [27] to convert the ${{P}_{i}}$ into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes are able to fit free-form objects efficiently and are well-suited for complex topological structures. The meshing process of the point cloud is illustrated in Figure 3. Subsequently, the tetrahedral mesh data is processed by the mesh processor $\mathrm{T}$ to be mapped into a format suitable for input to encoder ${{\Phi}_{\theta}}$. Where the mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, serving as a bridge between the mesh data and encoder ${{\Phi}_{\theta}}$. The encoder ${{\Phi}_{\theta}}$ maps the processed data to the $m$-dimensional latent space to obtain the latent vector ${z_{phys}}$. ${z_{phys}}$ is concatenated with the deformation parameters $d$ as the input $z_{phy}^{d}$ to decoder ${{\Psi}_{phys}}$, enabling ${{\Psi}_{phys}}$ to predict the displacement field $\hat{u}$ of the object. Decoder ${{\Psi}_{phys}}$ is composed of 5-layer of MLP, outputting a feature vector of size $n\times 3$, where $n$ is the number of input points in the point cloud.
We set parameter values in $d$ based on the characteristics of the point cloud object to simulate its properties, including Lamé constants $\lambda$ and $\mu$, elastic modulus $E$, Poisson’s ratio $\nu$, and external forces $f$. These parameters together determine the deformation behavior of the object. At the same time, based on the configuration of deformation parameters $d$, we obtain the ground-truth displacement field $u$ using the finite element analysis method (FEM) [9]. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output adheres to the physical laws. Below, we provide a detailed description of these two loss functions.
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field ${u}$, enabling it to precisely predict local geometric changes. In calculating elastic deformation, the finite element method is a common approach to solving elastic dynamics models in contact problems. Given $n$ node positions $x$, a finite element mass matrix $M$, and a hyper-elastic deformation energy $E(x)$, the solution to the contact problem can be obtained using the extremization method of the action:
[…]
where $S(x)$ represents the action of the system over the time interval $\left[{0,T}\right]$, describing the total energy of the object. $\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}$ denotes the kinetic energy, $E(x)$ represents the potential energy (elastic deformation potential energy) and ${{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)$ corresponds to the external force action. ${{f}_{e}}$ are external forces and ${{f}_{d}}$ are dissipative frictional forces.
Figure 3: The point cloud is first converted into a rough mesh form through Delaunay triangulation. Then, we removed the larger tetrahedral meshes, further refining them into the final tetrahedral representation.
The above describes the general analysis of dynamic problems. To simulation and analyze the deformation of an object under static linear elastic forces within the spatial domain ${\Omega}$, we use FEM to establish the stress-strain equation and solve for the displacement field $u$:
$S\left(u\right)=\int_{\Omega}{\left(\frac{1}{2}\sigma\colon\varepsilon\left(u\right)-f\cdot u\right)}d\Omega$ (4)
where ${\Omega}$ represents the volume region of the object, and $S(u)$ computes the total energy of the object’s deformation. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $-f\cdot u$ represents the potential energy exerted by the external force, which is the work done by the external force under the displacement field $u$.
We determine the force application direction by computing the object’s inertia matrix [35] to ensure that the deformation direction better aligns with the object’s geometric distribution, avoiding potential unreasonable force application caused by manual selection. The eigenvectors of the inertia matrix reflect the primary distribution trends of the object in different directions, providing an optimal force application direction that conforms to its geometric structure, thereby ensuring a more reasonable force application. Finally, we design the following data-fidelity loss:
${{\rm{{\cal L}}}_{df}}=\frac{1}{{\left|{\rm{{\cal D}}}\right|M}}\sum\limits_{i=1}^{\left|{\rm{{\cal D}}}\right|}{\sum\limits_{l=1}^{M}{{{\left\|{{u_{il}}-{{\hat{u}}_{il}}}\right\|}_{2}}}}$ (5)
where ${u_{il}}$ and ${{\hat{u}}_{il}}$ represent the ground truth displacement field of the tetrahedral mesh and the displacement field predicted by the network, respectively. $M$ denotes the number of tetrahedral meshes contained in the point cloud ${P_{i}}$. $\left\|\cdot\right\|_{2}$ represents the ${l_{2}}$-norm.
Physics-informed Loss. Based on the network’s prediction of the displacement of each tetrahedral mesh vertex, we derived the distribution of external forces on the tetrahedron using Hooke’s law [3] in linear elasticity analysis and the equilibrium equations [10]. First, each tetrahedron is composed of four vertices, denoted as ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ (initial coordinates) and ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (coordinates after deformation). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$ learned by the network. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
$F=X^{\prime}\cdot{X^{-1}}$ (6)
After obtaining the deformation gradient, the strain tensor $\varepsilon$ of the tetrahedral mesh can be calculated using the following formula. $\varepsilon$ reflects the deformation information of each tetrahedron element, including both the stretching and shear components, as shown in the following equation:
$\varepsilon=\frac{1}{2}\left(F+{{F}^{T}}\right)-I$ (7)
where ${F}^{T}$ is the transpose matrix of the deformation gradient, and $I$ is the identity matrix.
In elasticity, stress $\sigma$ and strain $\varepsilon$ are related through constitutive relations. For linear elastic materials, Hooke’s law is usually used for description. In the three-dimensional cases, the stress tensor is calculated as:
$\sigma=\lambda Tr\left(\varepsilon\right)I+2\mu\varepsilon$ (8)
where $\lambda=\frac{E\nu}{\left(1+\nu\right)\left(1-2\nu\right)}$ and $\mu=\frac{E}{2\left(1+\nu\right)}$ is the Lamé constant, which is related to the elastic modulus $E$ and Poisson’s ratio $\nu$. These parameters, along with the external force $f$, determine how the object will deform.
In three-dimensional linear elasticity, the equilibrium equation is $\nabla\cdot\sigma+f=0$. Therefore, our physics-informed loss term is:
${{\rm{{\cal L}}}_{pi}}=\frac{1}{{\left|D\right|M}}\sum\limits_{i=1}^{\left|D\right|}{\sum\limits_{l=1}^{M}{{{\left\|{\nabla\cdot{\sigma_{il}}+{f_{il}}}\right\|}_{2}}}}$ (9)
where ${\nabla\cdot{\sigma_{il}}}$ represents the divergence of the stress tensor, which describes the force equilibrium within the mesh.
3.4 Overall Object
Finally, we integrate the loss functions ${{\mathcal{L}}_{imp}}$, ${{\mathcal{L}}_{df}}$, and ${{\mathcal{L}}_{pi}}$ during the training phase, where ${{\mathcal{L}}_{imp}}$ represents the implicit loss, ${{\mathcal{L}}_{df}}$ represents the data-fidelity loss, and ${{\mathcal{L}}_{pi}}$ represents the physics-informed loss. $a$ and $b$ are coefficient.
$\mathcal{L}_{all}={{\mathcal{L}}_{imp}}+a{{\mathcal{L}}_{df}}+b{{\mathcal{L}}_{pi}}$ (10)
4 Experiment
4.1 Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: cTree [37], Jigsaw3D [36], and Rotation [29] are based on pretext tasks. OcCo [44], GSPCon [13], and Tsai et al. [41] are based on the generative models. STRL [16], PoCCA [46], CLR-GAM [26], PointSmile [22], and ToThePoint [21] are based on contrast learning methods. CrossPoint [1], CrossNet [47], TCMSS [12], PointMCD [54], and Cosine Mixup [18] are based on cross-modal methods. Additionally, for clarity in the tables below, we use the abbreviations Sup, GM, PT, CL, and CM to represent supervised learning, generative models, pretext tasks, contrastive learning, and cross-modal methods, respectively.
Datasets. We use the ShapeNet [5] dataset as the pre-training dataset, which consists of 57,448 synthetic models from 55 categories. For downstream tasks, we evaluate our method on four different benchmark datasets for object classification and object segmentation. Object Classification: the synthetic dataset ModelNet40 [48] and the real-world dataset ScanObjectNN [42]. Object Segmentation: the synthetic dataset ShapeNetPart [51] and the real-world dataset S3DIS [2].
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [16] and use the backbone networks of PointNet [30] and DGCNN [45] as the encoder ${{\Phi}_{\theta}}$. We use Adam [17] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [25], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${{\Phi}_{\theta}}$. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [40]. We also evaluate the few-shot results on two classification datasets. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results.
Table 1: Comparison of linear classification accuracy based on ModelNet40 dataset.
Method Category PointNet Acc DGCNN Acc
PointNet [30] Sup 89.2 --
DGCNN [45] -- 92.2
Jigsaw3D [36] PT 87.3 90.6
Rotation [29] 88.6 90.8
STRL [16] CL 88.3 90.9
ToThePoint [21] $85.6$ 89.2
CLR-GAM [26] 88.9 91.3
PoCCA [46] 89.4 91.4
PointSmile [22] 90.0 91.8
CrossPoint [1] CM 89.1 91.2
PointMCD [54] 90.0 --
TCMSS [12] 87.7 91.4
Cosine Mixup [18] 88.7 91.7
CrossNet [47] 89.5 91.5
OcCo [44] GM 88.7 89.2
GSPCon [13] 89.8 91.5
Tsai et al. [41] 90.1 92.0
Ours 90.3 92.2
4.2 Object Classification
[…]
Table 2: Comparison of linear classification accuracy based on ScanObjectNN dataset.
Method Category PointNet Acc DGCNN Acc
PointNet [30] Sup 73.3 --
DGCNN [45] -- 82.8
Jigsaw3D [36] PT 55.2 59.5
STRL [16] CL 74.2 77.9
ToThePoint [21] 74.7 81.9
CLR-GAM [26] 75.7 82.1
PoCCA [46] 75.6 82.2
Pointsmile [22] 75.8 82.8
CrossPoint [1] CM 75.6 81.7
TCMSS [12] 75.1 84.0
Cosine Mixup [54] 72.5 84.7
CrossNet [47] 76.8 83.9
OcCo [44] GM 69.5 78.3
GSPCon [13] 75.8 83.9
Tsai et al. [41] 70.6 74.8
Ours 77.3 86.4
[…]
Table 3: Few-shot object classification results on ModelNet40 dataset. We report mean and standard error over 10 runs.
Method 5-way 10-way
10-shot 20-shot 10-shot 20-shot
Jigsaw3D [36] 34.3$\pm$1.3 42.2$\pm$3.5 26.0$\pm$2.4 29.9$\pm$2.6
cTree [37] 60.0$\pm$2.8 65.7$\pm$2.6 48.5$\pm$1.8 53.0$\pm$1.3
OcCo [44] 90.6$\pm$2.8 92.5$\pm$1.9 82.9$\pm$1.3 86.5$\pm$2.2
GSPCon [13] 93.6$\pm$2.9 95.3$\pm$2.5 86.1$\pm$4.7 89.9$\pm$3.7
ToThePoint [21] 92.7$\pm$4.8 95.1$\pm$3.0 87.9$\pm$4.3 91.1$\pm$3.6
Cosine Mixup [18] 92.9$\pm$3.9 95.3$\pm$2.4 87.2$\pm$2.8 90.6$\pm$3.1
CrossPoint [1] 92.5$\pm$3.0 94.9$\pm$2.1 83.6$\pm$5.3 87.9$\pm$4.2
CLR-GAM [26] 93.7$\pm$1.2 96.0$\pm$2.6 87.9$\pm$2.7 91.1$\pm$1.9
TCMSS [12] 93.1$\pm$4.2 95.4$\pm$5.3 84.5$\pm$5.1 90.5$\pm$3.5
PoCCA [46] 93.5$\pm$3.7 92.1$\pm$3.6 88.1$\pm$5.3 90.9$\pm$4.0
Ours 95.7$\pm$2.6 97.8$\pm$1.9 90.4$\pm$2.7 94.8$\pm$3.0
Table 4: Few-shot object classification results on ScanobjectNN dataset. We report mean and standard error over 10 runs.
Method 5-way 10-way
10-shot 20-shot 10-shot 20-shot
Jigsaw3D [36] 65.2$\pm$3.8 72.2$\pm$2.7 45.6$\pm$3.1 48.2$\pm$2.8
cTree [37] 68.4$\pm$3.4 71.6$\pm$2.9 42.4$\pm$2.7 43.0$\pm$3.0
OcCo [44] 72.4$\pm$1.4 77.2$\pm$1.4 57.0$\pm$1.3 61.6$\pm$1.2
GSPCon [13] 78.1$\pm$3.9 81.5$\pm$3.5 64.8$\pm$3.8 73.5$\pm$2.1
ToThePoint [21] 78.1$\pm$7.9 83.8$\pm$6.1 66.7$\pm$5.4 74.2$\pm$5.0
Cosine Mixup [18] 79.6$\pm$4.6 86.4$\pm$5.0 68.5$\pm$3.2 77.0$\pm$2.3
CrossPoint [1] 74.8$\pm$1.5 79.0$\pm$1.2 62.9$\pm$1.7 73.9$\pm$2.2
CLR-GAM [26] 80.6$\pm$1.9 86.3$\pm$2.3 67.2$\pm$1.5 76.4$\pm$2.7
TCMSS [12] 79.8$\pm$6.9 83.8$\pm$6.4 67.2$\pm$3.7 74.8$\pm$4.0
PoCCA [46] 79.9$\pm$4.7 83.5$\pm$4.2 66.0$\pm$3.2 75.1$\pm$2.7
Ours 86.4$\pm$3.7 90.3$\pm$3.8 76.1$\pm$2.1 82.5$\pm$3.3
Table 5: Performance comparison of segmentation under ShapeNetPart dataset and S3DIS dataset (DGCNN as backbone).
Method Category ShapeNetPart S3DIS
DGCNN [45] Sup 85.1 54.9
Jigsaw3D [36] PT 84.3 55.6
STRL [16] CL 85.1 --
ToThePoint [21] 85.5 --
Pointsmile [22] 85.4 58.9
CLR-GAM [26] 85.5 --
PoCCA [46] 85.8 --
CrossPoint [1] CM 85.3 58.4
PointMCD [54] 85.1 --
TCMSS [12] 85.8 58.8
CrossNet [47] 85.5 --
OcCo [44] GM 85.0 58.0
GSPCon [13] 85.7 --
Tsai et al. [41] 85.1 --
Ours 86.0 58.9
Figure 4: Visualisation of part segmentation results based on DGCNN backbone. Different colours indicate different parts. Top row: ground-truth; bottom row: The results of our projections.
Figure 5: Self-supervised training of the DGCNN network backbone is performed on the ShapeNet dataset, and t-SNE visualization of features is tested on the ModelNet10 dataset. (a) represents the result of using only physical loss. (b) represents the result of using only the implicit field loss. (c) represents our method.
[…]
4.3 Object Segmentation
We transfer the pre-trained DGCNN backbone model to the segmentation task and conduct experimental analysis using the ShapeNetPart and S3DIS datasets, fine-tuning the model in an end-to-end manner on both datasets. We report the mean IoU (Intersection-over-Union) metric in Table 5. Specifically, we perform 6-fold cross-validation on the 6 regions of the S3DIS dataset to obtain the final segmentation results. It can be seen that our method performs well on both datasets, demonstrating that the incorporation of physical information enhances the model’s segmentation capability. Figure4 shows the segmentation results on the ShapeNetPart dataset. Our method effectively segments the fine-grained details of the parts, demonstrating superior part segmentation performance.
4.4 Ablation Study
We qualitatively analyzed the effect of adding physical information on the whole detection performance. As shown in Figure 5, t-SNE [43] draws a visualization of the inputfeatures based on the ModelNet10 dataset. It can be observed that the model guided solely by physical information (as shown in (a)) is capable of learning feature representations of objects, although with relatively coarse granularity and indistinct boundaries between different object features. Furthermore, when the physics-based model is incorporated as a constraint within the data-driven pipeline (as shown in (c)), the feature learning performance significantly improves. Compared to purely data-driven methods (as shown in (b)), this hybrid methodology achieves sharper distinctions between object features and exhibits enhanced learning capabilities for characterizing objects. For some classes (e.g., desk, sofa) have imprecise class boundaries, our method still does a better job of distinguishing them.
In addition, we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Similar to the classification experiments, we performed pre-training for each part separately, followed by training an SVM classifier using the extracted point cloud features, as summarized in Table 6. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
Table 6: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 and ScanObjectNN datasets.
Datasets ${{\cal L}_{df}}+{{\cal L}_{pi}}$ ${{\cal L}_{imp}}$ ${{\cal L}_{df}}+{{\cal L}_{pi}}+{{\cal L}_{imp}}$
ModelNet40 85.1 90.5 92.2
ScanObjectNN 78.3 84.1 86.4
5 Conclusion
This paper proposes a physics-driven self-supervised learning method that learns the whole features of an object using implicit fields while establishing a deformation mapping from local to whole through displacement prediction. This mapping forms a closed-loop feedback mechanism that deeply integrates local and whole geometric features. Experimental results demonstrate that our method outperforms existing methods in downstream tasks such as object classification, few-shot learning, and segmentation, validating its effectiveness and robustness. This work introduces a new physically-guided paradigm for point cloud representation learning, significantly enhancing the ability to learn point cloud features.