### v2

# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

Abstract
Existing point cloud representation learning methods primarily rely on data-driven strategies to extract geometric information from large amounts of scattered data. However, most methods focus solely on the spatial distribution features of point clouds while overlooking the relationship between local information and the whole structure, which limits the accuracy of point cloud representation. Local information reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric features. Therefore, the appropriate introduction of physics-driven mechanism can effectively compensate for the limitations of data-driven methods in structural modeling and significantly enhance the generalization and interpretability of point cloud representations in downstream tasks such as understanding and recognition. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method to learn fine-grained features in point clouds and model the structural relationship between local regions and the whole shape. Specifically, we design a dual-task encoder-decoder framework that combines the geometric modeling capability of data-driven implicit fields with physics-driven elastic deformation. Through the integration of physics-based loss functions, the framework is guided to predict localized deformation and explicitly capture the correspondence between local structural changes and whole shape variations. Experimental results show that our method outperforms existing approaches in object classification and segmentation, demonstrating its effectiveness.
Index Terms:
Point Cloud Representation, Physics-driven, Implicit Field, Self-supervised Learning.
I Introduction
Point clouds can accurately describe the geometric shape and spatial distribution of objects. They are widely used in tasks such as 3D shape analysis [1], object detection [2], and segmentation [3, 4]. The success of these tasks depends on the accurate modeling of point cloud features, while the irregularity and sparsity of point clouds make learning their effective representations a challenging but essential task.
Figure 1: (a) Visualization of deformation and force distribution under an external force $f$ for Lamp A and B. Colors represent force magnitude, with red circles indicating regions of significant deformation. (b) Clustering results with and without Physics-Informed (PI) modeling.
Existing methods primarily rely on data-driven approaches to capture geometric distributions of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [5], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. PointNet++ [6] builds upon this by introducing a hierarchical structure to capture local features, allowing for better capture of structural information in local regions. DGCNN [7] utilizes dynamic graph convolutional networks to capture local geometric structural features. In unsupervised point cloud representation learning, generative methods [8, 9, 10, 11] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [12], generative adversarial networks (GAN) [13], or autoregressive (AR) [10]. Discriminative methods [14, 15, 16, 17] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [18, 19] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
Specifically, existing methods often expand local information to the global level through multi-scale operations (e.g., the hierarchical structure of PointNet++) or max-pooling operations (e.g., global feature aggregation in PointNet). This process of aggregating information solely in the feature space does not explicitly model and capture the interrelationship between local and whole structures. For example, DGCNN captures local geometric structures using dynamic graph convolutional networks, but its layer-by-layer expansion of local information remains a passive, lacking explicit modeling of the relationship between part and whole structural features.
In nature, the shape of objects is often governed by the forces acting upon them. When an object is subjected to external forces, slight deformations first occur in its local regions. These deformations propagate through structural connections, causing global shape changes. This transition from “local stress” to “global response” reveals the inherent coupling relationships among the parts of an object. Meanwhile, recent research [20] has shown that physical modeling not only provides interpretability and inductive capability, but can also be embedded into deep networks as structural constraints to enhance the performance of data-driven methods on downstream detection tasks. As shown in Figure 1, incorporating the physics-driven approach enables more effective discrimination of fine-grained features both intra-class and inter-class. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method, which captures the deformation effects between local and whole shapes by introducing the force propagation mechanism from elasticity theory. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of object from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, simulating the force propagation relationship between local and whole structures. The main contributions of our method are as follows:
$\bullet$ We propose a novel self-supervised learning method that explicitly captures the relationship between local features and whole structure in point clouds by combining physical laws with geometric modeling. This is also the first self-supervised learning method that leverages the mechanism of force propagation to model the relationship between local and whole structures of objects.
$\bullet$ We designed two key loss functions to learn the physical characteristics of objects: a data-fidelity loss and a physics-informed loss. These two loss functions work synergistically during the optimization process to enforce physical consistency in the point cloud.
$\bullet$ We demonstrate the effectiveness of our method through extensive evaluations on multiple downstream tasks, including object classification and segmentation. Additionally, the incorporation of physical information enhances the understanding of 3D point clouds.
II Related Work
II-A Point Cloud Representation Learning
In recent years, the advancement of deep learning techniques has enabled direct utilization of point cloud data for learning intrinsic feature representations of shapes. Early methods, such as PointNet [5], used shared MLPs and max pooling operations to learn global features of point clouds. Subsequently, PointNet++ [6] introduced a hierarchical structure and local feature learning mechanisms to address the issue of PointNet’s insensitivity to local features in point clouds. DGCNN [7] further improves local feature extraction by dynamically constructing graphs based on point neighborhoods. Similarly, recent studies [21, 22, 23] have focused on learning local feature information from point clouds, aiming to extract as many semantic features as possible. However, the methods often treat point clouds as static geometric distributions, overlooking the explicit relationship between local and whole structures. This limitation, especially in the case of object deformation, makes it difficult to capture the impact of local changes on the whole structure. Against this backdrop, implicit field modeling [24, 25] has been gradually introduced into point cloud learning, representing the geometric structure of point clouds through continuous functions. This representation inherently provides spatial continuity and resolution adaptivity, making it well-suited for capturing the continuous relationship between local and whole structures. It also offers a stronger geometric foundation for subsequent unsupervised modeling.
II-B Unsupervised Point Cloud Representation Learning
Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [26, 8, 27, 28, 11, 29]. For example, HC-VAE [27] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [28] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [9, 10, 8, 30] has also become a prominent research focus. MaskSurf [9] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [8] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process.
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [31, 32, 14, 33, 34, 35]. Jigsaw3D [31] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [36] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [15, 37, 14] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [14] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [15] is an extension of the method [38] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [17] guides network learning by maximizing the mutual information between two point clouds. In addition, some researchers have focused on incorporating cross-model [39, 40, 41, 16, 42, 43, 44] methods into discriminative methods. However, these methods often overlook the influence of local structures on whole shape responses and typically rely on large numbers of positive-negative sample pairs and complex strategies, resulting in a learning process that lacks physical interpretability.
Figure 2: The overall architecture of the proposed method is illustrated. It consists of two modules: the implicit feature learning module, which represents the point cloud as a continuous surface, helping the network better learn the whole shape features of the object; and the physics information awareness module, which learns the relationship between local features and whole geometry through the elastic deformation of the object, enhancing the model’s ability to learn local details.
In this work, we incorporate a physics-driven mechanism into the data-driven method that explicitly models deformation propagation from local forces to whole structures, enabling physically consistent and interpretable point cloud understanding. This also leads to more robust and generalizable features for downstream tasks.
III Method
The framework of our method (Figure 2) consists of two main modules. The first module is the implicit feature learning module, where the point cloud is input into an autoencoder to learn its implicit features. The implicit field represents the point cloud as a continuous surface, and by leveraging the distance function from a query point to the nearest surface, it can uniformly handle objects of arbitrary shapes, helping the network better learn the whole shape features of the object. The second module is the physical information awareness module, which captures local deformation behaviors under external forces by constructing tetrahedral meshes and introducing deformation mechanisms from elasticity theory. This module not only learns the physical characteristics of local responses, but also embeds local deformation features into a unified geometric representation space via a shared encoder, thereby enabling mutual enhancement between whole shape representation and local physical perception. The proposed dual-task self-supervised framework combines the interpretability of physics-based modeling with the strong representational capacity of data-driven methods, enabling more effective learning of the geometric representations of real-world objects.
III-A Implicit Feature Learning
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${\Phi}_{\theta}$} and decoder ${\Psi}_{\eta}$} to learn the implicit features of the point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$. The encoder ${\Phi}_{\theta}$} takes the $P_{i}$} as input and maps it to an $m$-dimensional latent vecoter $z_{\eta}$}, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$, where we use the network backbones of PointNet [5] or DGCNN [7] as ${\Phi}_{\theta}$} to learn the point cloud features, respectively. Subsequently, we concatenate latent vector $z_{\eta}$} with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Then, the decoder ${\Psi}_{\eta}$} takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud $P_{i}$}. The decoder ${\Psi}_{\eta}$} consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
Our training objective is to minimize the error between the distance from the query point $q$ to the implicit surface of the point cloud ${P}_{i}$, as predicted by the implicit function $\delta_{i}({z_{\eta}^{q}})$, and the ground truth distance ${\delta}_{gt}(q)$:
${\theta^{*}},{\eta^{*}}=\mathop{\arg\min}\limits_{\theta,\eta}\left({\delta_{i}(z_{\eta}^{q})),{\delta_{gt}}\left(q\right)}\right),$ (1)
where ${\delta}_{gt}(q)$ represents the nearest neighbor distance from the query point $q$ to the points on the point cloud ${P}_{i}$}, expressed using the unsigned distance function value [45]. Therefore, the implicit loss ${\cal L}_{imp}$ is defined as:
${{\cal L}_{imp}}=\frac{1}{{|\mathcal{D}|K}}\sum\limits_{i=1}^{|\mathcal{D}|}{\sum\limits_{q\in Q}^{K}{\left|{\left|{{\delta_{i}}(z_{\eta}^{q})}\right|-{\delta_{gt}(q)}}\right|}},$ (2)
where $K$ is the number of the query point $q$ randomly sampled within the unit space, and $Q$ is the set of query points.
III-B Physical Information Awareness
For a point cloud ${P}_{i}$}, we first use 3D Delaunay [46] to convert the ${P}_{i}$} into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes better fit free-form objects and support complex topological structures, providing the volumetric elements required for physical modeling, such as deformation gradients and stress tensor computation. Meanwhile, directly using point coordinates lacks structural information, making it difficult to establish the strain–stress relationships necessary for constructing the subsequent physics-informed loss ${\mathcal{L}}_{pi}$. The advantages of using tetrahedral structures over point-based representations will be further validated in the ablation studies. The meshing process of the point cloud is illustrated in Figure 3.
Figure 3: The point cloud is first converted into a rough mesh form through Delaunay triangulation. Then, we removed the larger tetrahedral meshes, further refining them into the final tetrahedral representation.
Subsequently, the tetrahedral mesh data is processed by the mesh processor $\mathrm{T}$, which encodes the local geometric and topological structure of each tetrahedron into a unified format suitable for input to the encoder ${\Phi}_{\theta}$}, where the mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, serving as a bridge between the mesh data and encoder ${\Phi}_{\theta}$}. The encoder ${\Phi}_{\theta}$} maps the processed data to the $m$-dimensional latent space to obtain the latent vector $z_{phys}$}. $z_{phys}$} is concatenated with the deformation parameters $d$ as the input $z_{phy}^{d}$ to decoder ${\Psi}_{phys}$}, enabling the model to learn local-to-whole deformation patterns under varying physical conditions during training. The $d$ includes standard physical parameters commonly used in static linear elasticity theory, such as Lamé constants $\lambda$ and $\mu$, elastic modulus $E$, Poisson’s ratio $\nu$, and external forces $f$. These parameters have clear physical meanings and can effectively characterize stress responses of typical elastic materials under small deformations. Considering the differences in point cloud density and actual deformation scale, we further apply an adaptive adjustment to the parameters in $d$ based on the size of the generated tetrahedral mesh and the initial deformation magnitude. This reduces reliance on manual configuration and enhances the robustness and generalization ability of the proposed method. Decoder ${\Psi}_{phys}$} is composed of 5-layer of MLP, outputting a feature vector of size $n\times 3$, where $n$ is the number of input points in the point cloud.
Based on the configuration of deformation parameters $d$, we obtain the ground-truth displacement field $u$ using the finite element analysis method (FEM) [47]. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output adheres to the physical laws. Below, we provide a detailed description of these two loss functions.
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field u{u}, enabling it to precisely predict local geometric changes. In calculating elastic deformation, the finite element method is a common approach to solving elastic dynamics models in contact problems. Given $n$ node positions $x$, a finite element mass matrix $M$, and a hyper-elastic deformation energy $E(x)$, the solution to the contact problem can be obtained using the extremization method of the action:
$S\left(x\right)=\int_{0}^{T}{\left(\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}-E\left(x\right)+{{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)\right)}dt,$ (3)
where $S(x)$ represents the action of the system over the time interval $\left[{0,T}\right]$, describing the total energy of the object. $\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}$ denotes the kinetic energy, $E(x)$ represents the potential energy (elastic deformation potential energy) and ${{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)$ corresponds to the external force action. ${f}_{e}$} are external forces and ${f}_{d}$} are dissipative frictional forces.
The above describes the general analysis of dynamic problems. To simulation and analyze the deformation of an object under static linear elastic forces within the spatial domain $\Omega$}, we use FEM to establish the stress-strain equation and solve for the displacement field $u$:
$S\left(u\right)=\int_{\Omega}{\left(\frac{1}{2}\sigma\colon\varepsilon\left(u\right)-f\cdot u\right)}d\Omega,$ (4)
where $\Omega$} represents the volume region of the object, and $S(u)$ computes the total energy of the object’s deformation. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $f\cdot u$ represents the potential energy associated with external body forces under the displacement field, which is the work done by the external force under the displacement field $u$.
We determine the force application direction by computing the object’s inertia matrix [48] to ensure that the deformation direction better aligns with the object’s geometric distribution, avoiding potential unreasonable force application caused by manual selection. The eigenvectors of the inertia matrix reflect the primary distribution trends of the object in different directions, providing an optimal force application direction that conforms to its geometric structure, thereby ensuring a more reasonable force application. Finally, we design the following data-fidelity loss:
${{\rm{{\cal L}}}_{df}}=\frac{1}{{\left|{\rm{{\cal D}}}\right|M}}\sum\limits_{i=1}^{\left|{\rm{{\cal D}}}\right|}{\sum\limits_{l=1}^{M}{{{\left\|{{u_{il}}-{{\hat{u}}_{il}}}\right\|}_{2}}}},$ (5)
where $u_{il}$} and ${\hat{u}}_{il}$} represent the ground truth displacement field of the tetrahedral mesh and the displacement field predicted by the network, respectively. $M$ denotes the number of tetrahedral meshes contained in the point cloud $P_{i}$}. $\left\|\cdot\right\|_{2}$ represents the $l_{2}$}-norm.
Physics-informed Loss. Based on the predicted displacements of tetrahedral vertices by the network, we compute the external force residuals by deriving stress tensors from predicted displacements using Hooke’s law [49] and evaluating their divergence based on the equilibrium equation [50]. This local deformation is propagated through the entire mesh via force equilibrium constraints, enforcing consistency across connected tetrahedra and shaping the whole displacement field $\hat{u}$. First, each tetrahedron is composed of four vertices, denoted as ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ (initial coordinates) and ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (coordinates after deformation). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$ learned by the network. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
$F=X^{\prime}\cdot{X^{-1}}$ (6)
After obtaining the deformation gradient, the strain tensor $\varepsilon$ of the tetrahedral mesh can be calculated using the following formula. $\varepsilon$ reflects the deformation information of each tetrahedron element, including both the stretching and shear components, as shown in the following equation:
$\varepsilon=\frac{1}{2}\left(F+{{F}^{T}}\right)-I,$ (7)
where ${F}^{T}$ is the transpose matrix of the deformation gradient, and $I$ is the identity matrix.
In elasticity, stress $\sigma$ and strain $\varepsilon$ are related through constitutive relations. For linear elastic materials, Hooke’s law is usually used for description. In the three-dimensional cases, the stress tensor is calculated as:
$\sigma=\lambda Tr\left(\varepsilon\right)I+2\mu\varepsilon,$ (8)
where $\lambda=\frac{E\nu}{\left(1+\nu\right)\left(1-2\nu\right)}$ and $\mu=\frac{E}{2\left(1+\nu\right)}$ is the Lamé constant, which is related to the elastic modulus $E$ and Poisson’s ratio $\nu$. $Tr\left(\varepsilon\right)$ is the trace of the strain tensor, representing the volumetric deformation. These parameters, along with the external force $f$ , determine how the object will deform.
In three-dimensional linear elasticity, the equilibrium equation is $\nabla\cdot\sigma+f=0$. Therefore, our physics-informed loss term is:
${{\rm{{\cal L}}}_{pi}}=\frac{1}{{\left|D\right|M}}\sum\limits_{i=1}^{\left|D\right|}{\sum\limits_{l=1}^{M}{{{\left\|{\nabla\cdot{\sigma_{il}}+{f_{il}}}\right\|}_{2}}}},$ (9)
where $\nabla\cdot{\sigma_{il}}$} represents the divergence of the stress tensor, which describes the force equilibrium within the mesh. $f_{il}$ denotes the ground-truth external force at the $l$-th tetrahedron of the $i$-th sample. This loss term evaluates whether each tetrahedral element satisfies physical equilibrium, thereby serving as a physics-based constraint to guide the network toward generating reasonable predictions that comply with mechanical laws.
III-C Overall Object
Finally, we integrate the loss functions ${\mathcal{L}}_{imp}$}, ${\mathcal{L}}_{df}$}, and ${\mathcal{L}}_{pi}$} during the training phase, where ${\mathcal{L}}_{imp}$} represents the implicit loss, ${\mathcal{L}}_{df}$} represents the data-fidelity loss, and ${\mathcal{L}}_{pi}$} represents the physics-informed loss. $a$ and $b$ are coefficient.
$\mathcal{L}_{all}={{\mathcal{L}}_{imp}}+a{{\mathcal{L}}_{df}}+b{{\mathcal{L}}_{pi}}$ (10)
IV Experiment
IV-A Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: Jigsaw3D [31] and Rotation [36] are based on pretext tasks. CPG[51], PointMoment[35], PointSmile [17], PoCCA [33], CLR-GAM [34], ToThePoint [37], and STRL [15] are based on contrast learning methods. TCMSS [39], HyperIPC[40], MCIB[41], CrossNet [16], Cosine Mixup [42], PointMCD [43], and CrossPoint [44] are based on cross-modal methods. GDANet[52], GSPCon [11], Tsai et al. [29], and OcCo [26] are based on the generative models. Additionally, for clarity in the tables below, we use the abbreviations Sup, PT, CL, CM, and GM to represent supervised learning, pretext tasks, contrastive learning, cross-modal methods, and generative models, respectively.
Datasets. We use the ShapeNet [53] dataset as the pre-training dataset, which consists of 57,448 synthetic models from 55 categories. For downstream tasks, we evaluate our method on four different benchmark datasets for object classification and object segmentation. Object Classification: the synthetic dataset ModelNet40 [54] and the real-world dataset ScanObjectNN [55]. Object Segmentation: the synthetic dataset ShapeNetPart [56] and the real-world dataset S3DIS [57].
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [15] and use the backbone networks of PointNet [5] and DGCNN [7] as the encoder ${\Phi}_{\theta}$}. We use Adam [58] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [59], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${\Phi}_{\theta}$}. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [60]. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results.
TABLE I: Linear classification accuracy (%) on ModelNet40 and ScanObjectNN. PN and DG denote PointNet and DGCNN backbones respectively.
Method Category ModelNet40 ScanObjectNN
PN DG PN DG
PointNet [5] Sup 89.2 - 73.3 -
DGCNN [7] - 92.2 - 82.8
Jigsaw3D [31] PT 87.3 90.6 55.2 59.5
Rotation [36] 88.6 90.8 - -
STRL [15] CL 88.3 90.9 74.2 77.9
ToThePoint [37] 85.6 89.2 74.7 81.9
CLR-GAM [34] 88.9 91.3 75.7 82.1
PoCCA [33] 89.4 91.4 75.6 82.2
PointSmile [17] 90.0 91.8 75.8 82.8
PointMoment [35] 88.8 90.9 79.0 80.5
CrossPoint [44] CM 89.1 91.2 75.6 81.7
Cosine Mixup [42] 88.7 91.7 72.5 84.7
CrossNet [16] 89.5 91.5 76.8 83.9
HyperIPC [40] - 91.8 - 84.5
TCMSS [39] 87.7 91.4 75.1 84.0
MCIB [41] - 91.6 76.6 85.4
OcCo [26] GM 88.7 89.2 69.5 78.3
Tsai et al. [29] 90.1 92.0 70.6 74.8
GSPCon [11] 89.8 91.5 75.8 83.9
Ours 90.3 92.2 79.5 86.4
Figure 4: We use t-SNE to visualise features extracted from PointNet and DGCNN on the ModelNet10 test set. Different color represents different categories.
IV-B Object Classification
[…]
TABLE II: Fine-tuning accuracy (%) of shape classification on the ModelNet40 dataset and the ScanObjectNN dataset. The self-supervised pre-trained model was used as the initial weights for the supervised learning method. PN and DG denote PointNet and DGCNN.
Method Category ModelNet40 ScanObjectNN
PN DG PN DG
PointNet [5] Sup 89.2 - 73.3 -
DGCNN [7] - 92.2 - 82.8
Jigsaw3D [31] PT 89.6 92.4 76.5 82.7
STRL [15] CL - 93.1 - -
PoCCA [33] 90.2 93.2 80.3 84.1
PointSmile [17] 90.7 93.0 - -
CPG [51] - 93.6 - -
PointMCD [43] CM 91.1 93.7 - -
CrossNet [16] - 93.4 - -
OcCo [26] GM 90.1 93.1 80.0 83.9
GDANet [52] 90.1 93.3 69.0 79.4
Ours 90.6 93.9 80.9 88.0
We evaluate the trained pre-trained weights on the ModelNet10 dataset by extracting features for each sample and performing t-SNE [61] clustering visualization, as shown in Figure 4. The left and right sides of the figure visualise the PointNet and DGCNN model features, respectively. It can be observed that our pre-trained model effectively distinguishes most samples according to their categories, resulting in clear classification boundaries.
TABLE III: Performance comparison of segmentation under ShapeNetPart dataset and S3DIS dataset (DGCNN as backbone).
Method Category ShapeNetPart S3DIS
DGCNN [7] Sup 85.1 54.9
Jigsaw3D [31] PT 84.3 55.6
STRL [15] CL 85.1 -
ToThePoint [37] 85.5 -
Pointsmile [17] 85.4 58.9
CLR-GAM [34] 85.5 -
PoCCA [33] 85.8 -
PointMoment [35] - 46.9
CrossPoint [44] CM 85.3 58.4
PointMCD [43] 85.1 -
TCMSS [39] 85.8 58.8
CrossNet [16] 85.5 -
MCIB [41] 86.2 -
OcCo [26] GM 85.0 58.0
GSPCon [11] 85.7 -
Tsai et al. [29] 85.1 -
GSPCon [11] 85.7 -
GDANet [52] - 59.0
Ours 86.0 59.0
[…]
Figure 5: Semantic segmentation results on S3DIS (DGCNN as backbone). Different colors indicate different objects. From left to right are the top view of the ground truth, the predicted top view, the front view, and the side view.
Figure 6: Visualisation of part segmentation results based on DGCNN backbone. Different colours indicate different parts. Top row: ground-truth; bottom row: The results of our projections.
IV-C Object Segmentation
To further validate the effectiveness of the physical mechanism in point cloud representation learning, we evaluate our method on object segmentation task. This task requires the model not only to understand the global geometric structure, but also to accurately capture the relative relationships between individual parts. We transfer the pre-trained DGCNN backbone model to the segmentation task and conduct experimental analysis using the ShapeNetPart and S3DIS datasets, fine-tuning the model in an end-to-end manner on both datasets. We report the mean IoU (Intersection-over-Union) metric in Table III. Specifically, we perform 6-fold cross-validation on the 6 regions of the S3DIS dataset to obtain the final segmentation results. It can be seen that our method performs well on both datasets, demonstrating that the incorporation of physical information enhances the model’s segmentation capability. Figure 5 and Figure 6 show the segmentation visualizations on the S3DIS and the ShapeNetPart dataset dataset, respectively. It can be observed that by introducing a force-driven local-to-global modeling mechanism, our method exhibits strong generalization ability across different fine-grained point cloud segmentation tasks, demonstrating a solid understanding of local details as well as the geometric relationships and structural layouts in complex indoor scenes.
Figure 7: Visualization of elastic deformation under external forces across object categories. The first row shows the original shapes of the objects, while the second row presents the deformation results predicted by the network. The red arrows indicate the direction of the applied force, while the green dashed circles highlight regions exhibiting significant deformation.
TABLE IV: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 and ScanObjectNN datasets.
Datasets ${{\cal L}_{df}}+{{\cal L}_{pi}}$ ${\cal L}_{imp}$} ${{\cal L}_{df}}+{{\cal L}_{pi}}+{{\cal L}_{imp}}$
ModelNet40 85.1 90.5 92.2
ScanObjectNN 78.3 84.1 86.4
TABLE V: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 dataset.
Test point cloud tetrahedral ${\cal L}_{df}$} ${\cal L}_{pi}$} accuracy
1 91.7
2 91.3
3 90.9
4 92.2
IV-D Ablation Study
Module Analysis. we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Similar to the classification experiments, we performed pre-training for each part separately, followed by training an SVM classifier using the extracted point cloud features, as summarized in Table IV. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
[…]
TABLE VI: The time and memory required for tetrahedral mesh generation and deformation calculations.
Type time/s memory/MB
[…]
Deformation visualization. Figure 7 presents the predicted geometric deformations of various object categories under downward external force. The top row shows the original shapes in the undeformed state, while the bottom row illustrates the network-predicted elastic deformations. It can be observed that significant local deformations occur in structurally critical regions, such as the handle of a cup, the brim edge of a hat, the grip of a pistol, the joint of a lamp arm, and the load-bearing surface of a chair. These areas are typically stress-sensitive zones where force transmission leads to stress concentration. In the figure, these regions with pronounced deformation effects are highlighted using green dashed circles. Importantly, these local deformations do not occur in isolation but emerge progressively under the influence of global force propagation through the structure, reflecting the network’s ability to model the physical coupling between local parts and the global shape. The experimental results validate the effectiveness and soundness of the proposed physics-driven modeling framework in simulating the deformation process from localized responses to holistic structural reconstruction, demonstrating its capacity to capture the internal force transmission patterns within complex objects.
V Conclusion
This paper proposes a physics-driven self-supervised learning method for point cloud representation, which effectively models structural responses under force propagation by introducing a local-whole deformation mechanism derived from elasticity theory. The framework consists of two training modules: one models the whole geometric structure using an implicit field, enabling expressive representation for arbitrary shapes; the other captures elastic responses induced by external forces based on tetrahedral mesh construction and constrains deformation behavior through a physics loss. Both branches share encoder weights, establishing a closed-loop mapping between local force-induced deformations and global shape changes, thereby deeply integrating local details with global structural features. Experimental results demonstrate that the proposed method achieves superior performance across multiple downstream tasks, including classification and segmentation, consistently outperforming existing state-of-the-art approaches. These results validate the effectiveness of incorporating physics-based modeling for enhancing the robustness and generalization capability of point cloud representations.
Nevertheless, the current method still has limitations. For instance, in complex non-rigid deformation scenarios, there remains room for improving the physical accuracy of tetrahedral mesh construction and deformation simulation. Additionally, how to further integrate force propagation modeling with object properties in real-world environments remains an important direction for future research. Subsequent work will focus on more efficient physical simulation techniques to advance point cloud understanding from purely geometric representations toward deeper integration with structural and physical attributes, enabling more interpretable and generalizable modeling of real-world objects.